OpenStudy (alina123):

Scientific Notation: 1. The body of a 150 lb person contains 2.3 * 10 to the -4 lb of copper.How much copper is contained in the bodies of 1200 such people? 2. The speed of light is appromiately 3 * 10 to the power of 8 m/s. How far does light travel in 6.0 * 10 to the first power seconds?

OpenStudy (anonymous):

1. The body of a 150 lb person contains 2.3 * 10 to the -4 lb of copper.How much copper is contained in the bodies of 1200 such people? \[2.3\times 10^{-4}\times1.2\times10^{3}\] \[(2.3\times1.2)\times (10^{-4+3})\] \[2.76\times10^{-1}\] .2. The speed of light is appromimately 3 * 10 to the power of 8 m/s. How far does light travel in 6.0 * 10 to the first power seconds? \[3\times10^{8}\times6.0\times10^{1}\] \[(3\times6)\times(10^{8+1})\] \[18\times10^9\] \[1.8\times10\times10^{9}\] \[1.8\times10^{10}\] Hope this helps!